# Extremal problems in BMO and VMO involving the Garsia norm

### v1

@@ Abstract.
where $\mathbb{D}$ is the open unit disk and $\mathcal{P}$ is the Poisson integral operator. The Garsia norm $\|f\|_{G}$ is then defined as $\sup_{z\in\mathbb{D}}\left(\Phi_{f}(z)\right)^{1/2}$, and the space ${\rm BMO}$ is formed by the functions $f\in L^{2}$ with $\|f\|_{G}<\infty$. If $\|f\|^{2}_{G}=\Phi_{f}(z_{0})$ for some point $z_{0}\in\mathbb{D}$, then $f$ is said to be a norm-attaining ${\rm BMO}$ function, written as $f\in{\rm BMO}_{\rm na}$. Note that ${\rm BMO}_{\rm na}$ contains ${\rm VMO}$, the space of functions with vanishing mean oscillation. We study, first, the functions $f$ in $L^{\infty}$ (as well as in $L^{\infty}\cap{\rm BMO}_{\rm na}$) with the property that $\|f\|_{G}=\|f\|_{\infty}$. The analytic case, where $L^{\infty}$ gets replaced by $H^{\infty}$, is discussed in more detail. Secondly, we prove that every function $f\in{\rm BMO}_{\rm na}$ with $\|f\|_{G}=1$ is an extreme point of ${\rm ball}\,({\rm BMO})$, the unit ball of ${\rm BMO}$ with respect to the Garsia norm. This implies that the extreme points of ${\rm ball}\,({\rm VMO})$ are precisely the unit-norm ${\rm VMO}$ functions. As another consequence, we arrive at an amusing “geometric” characterization of inner functions.
@@ 1. Introduction
so that $\mathcal{P}f$ is the harmonic extension of $f$ into ${\mathbb{D}}$. Finally, we recall that the Hardy space $H^{p}$, with $1\leq p\leq\infty$, is formed by the functions $f\in L^{p}$ for which $\mathcal{P}f$ is analytic on ${\mathbb{D}}$. Equivalently (see, e.g., [11, Chapter II]), a function $f$ from $L^{p}$ belongs to $H^{p}$ if and only if its Fourier coefficients
@@
We refer to [11, Chapter VI] or [14, Chapter X] for various alternative characterizations of (and equivalent norms on) ${\rm BMO}$. In addition, these sources offer a systematic presentation of the underlying theory, including the duality between $H^{1}$ and ${\rm BMOA}$, and a great deal more.
@@
where $c$ is a constant of modulus $1$, $B$ is a Blaschke product, and $S$ is a singular inner function; see [11, Chapter II]. To be more explicit, the factors involved are of the form
@@
will also interest us here. Such functions being bounded, we may equivalently speak of their membership in the set ${\rm VMO}\cap L^{\infty}(=:QC)$, known as the space of quasicontinuous functions, and in the corresponding analytic subspace
@@
For basic facts about $QC$ and $QA$, the reader may consult [16] or [11, Chapter IX]. Among other things, it is shown there that $QC$ (and hence $QA$) is a closed subalgebra of $L^{\infty}$, and that the inner functions in $QA$ are precisely the rational ones (i.e., finite Blaschke products).
@@
Our results are stated in Section 2 below. These chiefly concern the moduli of $G$-extremal functions on ${\mathbb{T}}$, the canonical factorization of such functions (in the analytic setting), and the geometry of the unit balls in the spaces ${\rm BMO}$ and ${\rm VMO}$, always endowed with the Garsia norm $\|\cdot\|_{G}$. The proofs are given in Sections 3 and 4. Finally, Section 5 (which can be viewed as a kind of addendum) contains an explicit—but somewhat technical—construction of a $G$-extremal outer function in $H^{\infty}$.
@@ 2. Statement of results
We begin by looking at the moduli of $G$-extremal functions. In particular, we show (see part (a) of the theorem below) that every nonnegative $L^{\infty}$ function on ${\mathbb{T}}$ is expressible as $|f|$ for some $G$-extremal function $f$ from a fairly small subalgebra of $L^{\infty}$, not just from $L^{\infty}$ itself. To be precise, the subalgebra in question is $H^{\infty}+C$, the linear hull of $H^{\infty}$ and $C:=C({\mathbb{T}})$. It is well known that $H^{\infty}+C$ is actually the smallest closed subalgebra of $L^{\infty}$ that properly contains $H^{\infty}$, while $QC$ is the largest self-adjoint subalgebra of $H^{\infty}+C$ (see [11, Chapter IX] or [16]).
@@
Of course, the necessity in (b) is trivial, since (2.2) characterizes the moduli of non-null $H^{\infty}$ functions (see, e.g. [11, Chapter II]). Also obvious is the sufficiency part in (c). Indeed, given a constant $c>0$, the function $f(z)=cz$ has modulus $c$ on ${\mathbb{T}}$, is $G$-extremal and belongs to ${\rm BMO}_{\rm na}\cap L^{\infty}$. This observation actually shows that statement (c) remains valid whenever the class ${\rm BMO}_{\rm na}\cap L^{\infty}$ is replaced by a subset thereof that contains the identity function $z$ (among such subsets we mention $QC$, $QA$ and ${\rm BMO}_{\rm na}\cap H^{\infty}$).
@@
It is a classical result (see, e.g., [11, Chapter II]) that the moduli of non-null $H^{p}$ functions, and hence the outer functions $\mathcal{O}_{\eta}$ in $H^{p}$, are characterized by the conditions $\eta\in L^{p}$ and $\log\eta\in L^{1}$; equivalently, the former condition should be coupled with (2.2). The ${\rm BMO}$ situation is subtler, and we refer to [8] for a description of the outer functions $\mathcal{O}_{\eta}$ lying in ${\rm BMOA}$. In [9], the membership of $\mathcal{O}_{\eta}$ in $QA$ was discussed (as part of a more general problem) and the appropriate criterion was furnished.
@@
Now consider an outer function $F\in QA$. It follows from Corollary 2.2, part (2), that $F$ is never $G$-extremal. At the same time, we already know that the $G$-extremality can be achieved by passing from $F$ to $IF$ if the inner factor $I$ is chosen appropriately. Our next result characterizes the inner functions $I$ that are eligible in this sense. The criterion that arises, as given in Theorem 2.5 below, might be contrasted with that in [5, Theorem 4], where the canonical factorization of a $QA$ function was treated. There, the outer factor was required to compensate for the oscillatory behavior of the inner one, while in the current setting the inner factor is needed to make the product oscillatory enough.
@@
with some constant $C_{f}\geq 0$. While $A^{\alpha}$ is usually normed by letting $\|f\|_{A^{\alpha}}$ be the best value of $C_{f}$ in (2.4), it is known (see [6, Lemma 1] or [7, Theorem 1]) that an equivalent norm is provided by the Garsia-type quantity
@@
It should be recalled that the extreme points of ${\rm ball}\,(L^{\infty})$ are precisely the unimodular functions, while those of ${\rm ball}\,(H^{\infty})$ are the functions $\varphi\in H^{\infty}$ with $\|\varphi\|_{\infty}=1$ for which (2.5) fails; see [4, Section V] or [13, Chapter 9]. Thus, our method of constructing a unit-norm non-extreme point of ${\rm ball}\,({\rm BMO})$ (resp., ${\rm ball}\,({\rm BMOA})$), as employed in Theorem 2.11 above, consists in multiplying a unit-norm non-extreme function $\varphi$ from ${\rm ball}\,(L^{\infty})$ (resp., ${\rm ball}\,(H^{\infty})$) by a suitable Blaschke product $B$; the latter is needed to make the resulting function $G$-extremal.
@@
Proof of Theorem 2.4. Assume first that $\|\eta\|_{\infty}=1$ (and hence $\|F\|_{\infty}=1$). In order that $F$ be $G$-extremal, it is necessary and sufficient that there exist a sequence $\{z_{n}\}\subset{\mathbb{D}}$ such that
@@ 5. Example: a $G$-extremal outer function
We finally mention, in connection with the above example, that our current construction is somewhat reminiscent of that in [8, Section 3], where a similar method was used to produce a non-${\rm BMOA}$ outer function with ${\rm BMO}$ modulus.

### v2

@@ Abstract.
where $\mathbb{D}$ is the open unit disk and $\mathcal{P}$ is the Poisson integral operator. The Garsia norm $\|f\|_{G}$ is then defined as $\sup_{z\in\mathbb{D}}\Phi_{f}(z)^{1/2}$, and the space ${\rm BMO}$ is formed by the functions $f\in L^{2}$ with $\|f\|_{G}<\infty$. If $\|f\|^{2}_{G}=\Phi_{f}(z_{0})$ for some point $z_{0}\in\mathbb{D}$, then $f$ is said to be a norm-attaining ${\rm BMO}$ function, written as $f\in{\rm BMO}_{\rm na}$. Note that ${\rm BMO}_{\rm na}$ contains ${\rm VMO}$, the space of functions with vanishing mean oscillation. We study, first, the functions $f$ in $L^{\infty}$ (as well as in $L^{\infty}\cap{\rm BMO}_{\rm na}$) with the property that $\|f\|_{G}=\|f\|_{\infty}$. The analytic case, where $L^{\infty}$ gets replaced by $H^{\infty}$, is discussed in more detail. Secondly, we prove that every function $f\in{\rm BMO}_{\rm na}$ with $\|f\|_{G}=1$ is an extreme point of ${\rm ball}\,({\rm BMO})$, the unit ball of ${\rm BMO}$ with respect to the Garsia norm. This implies that the extreme points of ${\rm ball}\,({\rm VMO})$ are precisely the unit-norm ${\rm VMO}$ functions. As another consequence, we arrive at an amusing “geometric” characterization of inner functions.
@@ 1. Introduction
so that $\mathcal{P}f$ is the harmonic extension of $f$ into ${\mathbb{D}}$. Finally, we recall that the Hardy space $H^{p}$, with $1\leq p\leq\infty$, is formed by the functions $f\in L^{p}$ for which $\mathcal{P}f$ is analytic on ${\mathbb{D}}$. Equivalently (see, e.g., [12, Chapter II]), a function $f$ from $L^{p}$ belongs to $H^{p}$ if and only if its Fourier coefficients
@@
We refer to [12, Chapter VI] or [15, Chapter X] for various alternative characterizations of (and equivalent norms on) ${\rm BMO}$. In addition, these sources offer a systematic presentation of the underlying theory, including the duality between $H^{1}$ and ${\rm BMOA}$, and a great deal more.
@@
where $c$ is a constant of modulus $1$, $B$ is a Blaschke product, and $S$ is a singular inner function; see [12, Chapter II]. To be more explicit, the factors involved are of the form
@@
will also interest us here. Such functions being bounded, we may equivalently speak of their membership in the space
$QC:={\rm VMO}\cap L^{\infty},$
whose elements are known as quasicontinuous functions, and in the corresponding analytic subspace
@@
For basic facts about $QC$ and $QA$, the reader may consult [17] or [12, Chapter IX]. Among other things, it is shown there that $QC$ (and hence $QA$) is a closed subalgebra of $L^{\infty}$, and that the inner functions in $QA$ are precisely the rational ones (i.e., finite Blaschke products).
@@
We also remark that the set ${\rm BMO}_{\rm na}$ is nonlinear, i.e., it does not form a vector space. Indeed, the singular inner function $S(z):=\exp\{(z+1)/(z-1)\}$ can be written as
$S=(1-z)S+zS$
and both terms on the right-hand side are in ${\rm BMO}_{\rm na}$, whereas $S$ is not.
Our results are stated in Section 2 below. These chiefly concern the moduli of $G$-extremal functions on ${\mathbb{T}}$, the canonical factorization of such functions (in the analytic setting), and the geometry of the unit balls in the spaces ${\rm BMO}$ and ${\rm VMO}$, always endowed with the Garsia norm $\|\cdot\|_{G}$. We also mention a new “geometric” characterization of inner functions that comes out in this context. The proofs are given in Sections 3 and 4. Finally, Section 5 (which can be viewed as a kind of addendum) contains an explicit—but somewhat technical—construction of a $G$-extremal outer function in $H^{\infty}$.
After this work had been completed, Kristian Seip kindly drew my attention to the recent paper [3], where a related extremal problem was treated in connection with Hankel operators. There, of concern were the functions $f\in H^{\infty}$ that satisfy
$\|f\|_{\infty}=\inf\left\{\|f-\overline{z}\overline{h}\|_{\infty}:\,h\in H^{\infty}\right\},$
a property reminiscent of (1.3). The class that arises turns out to have quite a bit in common with that of $G$-extremal functions lying in $H^{\infty}$. In particular, the similarity becomes apparent upon juxtaposing Theorem 3 from [3] with our Corollary 2.6, to be found in the next section.
@@ 2. Statement of results
We begin by looking at the moduli of $G$-extremal functions. Among other things, we show (see part (a) of the theorem below) that every nonnegative $L^{\infty}$ function on ${\mathbb{T}}$ is expressible as $|f|$ for some $G$-extremal function $f$ from a fairly small subalgebra of $L^{\infty}$, not just from $L^{\infty}$ itself. To be precise, the subalgebra in question is $H^{\infty}+C$, the linear hull of $H^{\infty}$ and $C:=C({\mathbb{T}})$. It is well known that $H^{\infty}+C$ is actually the smallest closed subalgebra of $L^{\infty}$ that contains $H^{\infty}$ properly, while $QC$ is the largest self-adjoint subalgebra of $H^{\infty}+C$; see [12, Chapter IX] or [17].
@@
Of course, the necessity in (b) is trivial, since (2.2) characterizes the moduli of non-null $H^{\infty}$ functions (see, e.g. [12, Chapter II]). Also obvious is the sufficiency part in (c). Indeed, given a constant $c>0$, the function $f(z)=cz$ has modulus $c$ on ${\mathbb{T}}$, is $G$-extremal and belongs to ${\rm BMO}_{\rm na}\cap L^{\infty}$. This observation actually shows that statement (c) remains valid whenever the class ${\rm BMO}_{\rm na}\cap L^{\infty}$ is replaced by a subset thereof that contains the identity function $z$ (among such subsets we mention $QC$, $QA$ and ${\rm BMO}_{\rm na}\cap H^{\infty}$).
@@
It is a classical result (see, e.g., [12, Chapter II]) that the moduli of non-null $H^{p}$ functions, and hence the outer functions $\mathcal{O}_{\eta}$ in $H^{p}$, are characterized by the conditions $\eta\in L^{p}$ and $\log\eta\in L^{1}$; equivalently, the former condition should be coupled with (2.2). The ${\rm BMO}$ situation is subtler, and we refer to [9] for a description of the outer functions $\mathcal{O}_{\eta}$ lying in ${\rm BMOA}$. In [10], the membership of $\mathcal{O}_{\eta}$ in $QA$ was discussed (as part of a more general problem) and the appropriate criterion was furnished.
@@
Now consider an outer function $F\in QA$. It follows from Corollary 2.2, part (2), that $F$ is never $G$-extremal. At the same time, we already know that the $G$-extremality can be achieved by passing from $F$ to $IF$ if the inner factor $I$ is chosen appropriately. Our next result characterizes the inner functions $I$ that are eligible in this sense. The criterion that arises, as given in Theorem 2.5 below, might be contrasted with that in [6, Theorem 4], where the canonical factorization of a $QA$ function was treated. There, the outer factor was required to compensate for the oscillatory behavior of the inner one, while in the current setting the inner factor is needed to make the product oscillatory enough.
@@
Remark. The above reasoning, and hence Corollary 2.6, remains valid if we merely assume that the outer function $F$ is in $H^{\infty}$ and $|F|$ is continuous on ${\mathbb{D}}\cup{\mathbb{T}}$. (These conditions are known to imply that $F\in QA$; see [18, Theorem 3].) The current version of the corollary is, however, good enough to produce numerous examples of analytic $G$-extremal functions with “nice” outer factors.
@@
with some constant $C_{f}\geq 0$. While $A^{\alpha}$ is usually normed by letting $\|f\|_{A^{\alpha}}$ be the best value of $C_{f}$ in (2.4), it is known (see [7, Lemma 1] or [8, Theorem 1]) that an equivalent norm is provided by the Garsia-type quantity
@@
It should be recalled that the extreme points of ${\rm ball}\,(L^{\infty})$ are precisely the unimodular functions, while those of ${\rm ball}\,(H^{\infty})$ are the functions $\varphi\in H^{\infty}$ with $\|\varphi\|_{\infty}=1$ for which (2.5) fails; see [5, Section V] or [14, Chapter 9]. Thus, our method of constructing a unit-norm non-extreme point of ${\rm ball}\,({\rm BMO})$ (resp., ${\rm ball}\,({\rm BMOA})$), as employed in Theorem 2.11 above, consists in multiplying a unit-norm non-extreme function $\varphi$ from ${\rm ball}\,(L^{\infty})$ (resp., ${\rm ball}\,(H^{\infty})$) by a suitable Blaschke product $B$; the latter is needed to make the resulting function $G$-extremal.
@@
Proof of Theorem 2.4. Let $F:=\mathcal{O}_{\eta}$. Assume first that $\|\eta\|_{\infty}=1$ (and hence $\|F\|_{\infty}=1$). In order that $F$ be $G$-extremal, it is necessary and sufficient that there exist a sequence $\{z_{n}\}\subset{\mathbb{D}}$ such that
@@ 5. Example: a $G$-extremal outer function
We finally mention, in connection with the above example, that our current construction is somewhat reminiscent of that in [9, Section 3], where a similar method was used to produce a non-${\rm BMOA}$ outer function with ${\rm BMO}$ modulus.